for each equation complete the solution
y= x/4 + 9 (16, )

Answer by stanbon(75887)

for each equation complete the solution
y= x/4 + 9 (16, )
-----
If x = 16, y = (16/4) + 9
y = 4 + 9
y = 13
===============
Cheers,
Stan H.
===============